OpenStudy (anonymous):

find the derivative (a) y=x^2f(x) (b) y=f(x)/x^2 (c) y= x^2/f(x) (d) y=(1+xf(x))/sqrtx

OpenStudy (anonymous):

f(x) is differentiable

OpenStudy (anonymous):

i will do the first one, the rest are similar by the product rule \((gf)'=g'f+f'g\) in our case \(g(x)=x^2, g'(x)=2x\) and \[f(x)=f(x), f'(x)=f'(x)\] is the best you can do, since you don't have an explicit description of \(f\) therefore the derivative of \[x^2f(x)\] is \[2xf(x)+x^2f'(x)\] by the product rule

OpenStudy (anonymous):

the rest require the quotient rule, but the method is the same

OpenStudy (anonymous):

is there any other way. the professor said to refrain from using the quotient rule if we could

OpenStudy (anonymous):

your professor is mistaken don't be scared of the quotient rule

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

whats the quotient rule again

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

never mind i got it